OpenStudy (anonymous):

the points A,B and C have coordinates <4,-1,5> and <8,0,6> and <5,-3,3> prove that angle BAC is a right angle.

OpenStudy (ajprincess):

Find BA nd AC. Then show that BA is perpendicular to AC. Can u?

OpenStudy (anonymous):

ive been trying but it seems to end up a mess

OpenStudy (anonymous):

Find slope of Ba and AC

OpenStudy (anonymous):

I think @ajprincess tried to say that

OpenStudy (anonymous):

got any drawing aids there to show me

OpenStudy (ajprincess):

Thanx for pointing out @sauravshakya

OpenStudy (anonymous):

Oh..... wait its a three dimensional figure

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Can we use vector

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

find BA and AC. OA=4i-1j+5k, OB=8i+6k and OC=5i-3j+3k BA=OA-OB AC=OC-OA then show that BA.AC=0 I mean show that their dot product is zero.

OpenStudy (ajprincess):

Can u do @sparksaflyen ?

OpenStudy (anonymous):

trying

OpenStudy (ajprincess):

oh k. if u need help plz ask.

OpenStudy (anonymous):

i got 5i-j+k, obviously doing something wrong here, if you got drawing aid please use

OpenStudy (anonymous):

make that just 5i-j

OpenStudy (ajprincess):

u got 4i-j for vat?

OpenStudy (ajprincess):

I meant to say 5i-j?

OpenStudy (anonymous):

obviously wrong

OpenStudy (anonymous):

maybe to many u-tube videos from educator is confusing me

OpenStudy (anonymous):

again if you got a drawing aid it would help

OpenStudy (ajprincess):

BA=4i-1j+5k-8i-6k =-4i-j-k AC=5i-3j+3k-4i+1j-5k =i-2j-2k BA.AC=(-4*1)+(-1*-2)+(-1*-2) =-4+2+2 =-4+4 =0 So BA is perpendicular to AC. So <BAC is a right angle.

OpenStudy (ajprincess):

Is this clear? Plz ask if it os nt clear. Will b back aftr some time.

OpenStudy (anonymous):

not clear

OpenStudy (anonymous):

why are you putting it the ijk form the coordinates are allready given i'm having trouble plotting these points

OpenStudy (anonymous):

example do i go x axis 4, y axis across -1, z axis up 5 and so on

OpenStudy (ajprincess):

r u askd to plot them?

OpenStudy (ajprincess):

I am really sorry. i feel it hard to plot them and show u in coordinate system.

OpenStudy (anonymous):

thats ok i know how hard it is to draw these things

OpenStudy (anonymous):

 this was the best i could do to draw them here :P

OpenStudy (anonymous):

never mind